# Automatic Hardware Pragma Insertion in High-Level Synthesis: A Non-Linear Programming Approach

### v2

@@ 4.2. A Formal Model for Latency
Our objective is to formulate a lower bound on the latency of a program after HLS. We therefore have put several restrictions: we assume the input program is a polyhedral program, that is the control-flow is statically analyzable; all loops can be recognized and their trip count computed; and all array / memory accesses can be exactly modeled at compile-time. No conditional can occur in the program. While our approach may generalize beyond this class, we limit here to these strict assumptions.
@@
We assume programs are made of affine loops, that are loops with statically computable control-flow, with loop bounds made only of intersection of affine expressions of surrounding loop iterators and program constants. We now assume loop bodies (i.e., statements surrounded by loops) have been translated to a list of statements, with at most a single operation (e.g., +,-,/,*) per statement. Operations are n-ary, that is they take $n\geq 0$ input scalar values as operand, and produce 0 or 1 output scalar value. A memory location can be loaded from (resp. stored to) an address stored in a scalar variable. This is often referred to as straight-line code . This normalization of the loop body facilitates the computation of live-in/live-out data for the code block, and the extraction of the computation graph. Note the region can be represented in Static Single Assignment form, to ensure different storage location for every assignment, facilitating the construction of the operation graph. In addition we require the input program to not contain useless operations which may be removed by the HLS toolchain e.g. by dead code elimination, as illustrated in Listing 4.
1int example(int y, int z){
2 int x;
3 S0: x = 12 + y; // dead-code elimination
4 S1: x = y + z;
5 return x;
6}
Listing 4: Example of Code Illustrating Dead Code Elimination: As Statement S1 writes to variable x, but the value of x assigned in Statement S0 is never used, the compiler will remove Statement S0.
@@ 4 S1: x = y + z;
  • No HLS optimization shall change the number of operations in the computation: strength reduction, common sub-expression elimination, etc. shall either first be performed in the input program before analysis, or not be performed by the HLS toolchain. The program also does not contain ”useless” operations that may be removed by the compiler.
@@
  • We assume resource (DSP) sharing across different operations executing at the same cycle is not possible.
An important term is $SL$, a latency lower bound for a region of straight-line code. To maintain a lower bound on latency by composition, we operate on a representation of (parts of the) program which is both schedule-independent and storage-independent: the operation graph, or CDAG (Elango et al., 2015). Indeed, a lower bound on this representation is necessarily valid under any schedule and storage eventually implemented and can be used to prove I/O lower bounds on programs (Elango et al., 2015) which is the directed acyclic graph with one node per operation in the code region, connecting all producer and consumer operations to build the operation graph. Then, we can easily compute the length of its critical path, which represents the minimal set of operations to execute serially.
@@
Given a straight-line code region $R$ made of a list $L$ of statements $S\in L$, the operation graph OG is the directed graph <$<\{N,V_{I},root,V_{O}\},E>$ such that $\forall S_{i}\in L,~{}N_{S_{i}}\in N$; and for every operation with output $o$ and inputs $\vec{i}$ in $L$ $\forall S_{i},\forall i_{k}\in\vec{i}_{S_{i}},e_{i_{k},o}\in E$, $\forall S_{i}:(o_{S_{i}},{\vec{i}}_{S_{i}})\in L,S_{j}:(o_{S_{j}},{\vec{i}}_{S_{j}})\in L$ with $S_{i}\neq S_{j}$ then we have $E_{S_{i},S_{j}}\in E$ iff $o_{Si}\cap{\vec{i}}_{S_{j}}\neq\emptyset$. For every input (resp. output) in $S_{i}$ which is not matched with an output (resp. input) of another $S_{j}$ in $L$, create a node $V_{val}\in V_{I}$ (resp. $V_{O}$) for this input (resp. output) value. If $dim({\vec{i}}_{S_{i}})=0$ then an edge $e_{root,S_{i}}$ is added to $E$.
From this representation, we can easily define key properties to subsequently estimate the latency and area of this code region, such as its span, or critical path.
@@
On the graph on the Figure 1 the critical paths from $A[i]$ to $c,\forall i\in$ $[\![0,7]\!]$, have the same length.
@@ 4.2.1. Latency Lower Bound
We can build a simple a lower bound on the latency of an operation graph:
@@
Given $R_{op}$ a count of available resources of type $op$, for each operation type, and $LO(op)$ the latency function for operation $op$, with $LO(op)\geq 1$. #$\#L(op)$ denotes the number of operations of type $op$ in $L$. We define $LO(\#OG_{cp}^{L})=\sum_{n\in cp}LO(n)$ the critical path weighted by latency of its operations. The minimal latency of a region $L$ is bounded by
$Lat_{R_{op}}^{L}\geq\max(LO(\#OG_{cp}^{L}),\max_{o\in op}(\lceil\#L(o)\times LO(o)/R_{o}\rceil))$
This theorem provides the building block to our analysis: if reasoning on a straight-line code region, without any loop, then building the operation graph for this region and reasoning on its critical path is sufficient to provide a latency lower bound.
1#pragma ACCEL unroll factor=uf
2L0: for (i = 0; i < N; i++)
@@ 10 S3: q[i]+=A[i][j]*p[j];
12}
Listing 5: Bicg code: $s=r\times A;q=A\times p$
For example, in Lst. 5, if we consider the sub-loop body composed of loops L2 (fully unrolled) and the statements $S2$ and $S3$ as straight-line code regions, we can calculate the critical paths for $S2$ and $S3$ as follows: For $S2$, the critical path is given by: $cp_{S2}=\max(L(+)+L(*),N\times(DSP_{+}+DSP_{*})/R_{o})$ with $DSP_{o}$ the number of DSP for the operation $o$. For $S3$, the critical path is determined by: $cp_{S3}=\max(L(+)\times\log(N)+L(*),N\times(DSP_{+}+DSP_{*})/R_{o})$, considering the possibility of a tree reduction. In this context, the critical path for the entire sub-loop body is the maximum of these two individual critical paths, expressed as $\max(cp_{S2},cp_{S3})$.
@@ 4.2.2. Loop Unrolling: partial unroll
Unrolling can be viewed as a two-step transformation: first strip-mine the loop by the unroll factor, then consider the inner loop obtained to be fully-unrolled. The latency of the resulting sub-program is determined by how the outer-loop generated will be implemented. We assume without additional explicit information this unrolled loop will execute in a non-pipelined, non-parallel fashion. Note this bound requires to build the operation graph for the whole loop body. This is straightforward for inner loops and/or fully unrolled loop nests, but impractical if the loop body contains other loops. We therefore define a weaker, but more practical, bound that enables composition:
@@
1#pragma ACCEL parallel UF=4
2L1: for (i = 0; i < 8; i++) {
3 S0: a[i] = b[i] * c[i];
4}
Listing 6: Example before partial unrolling
1// explicitly unroll with UF=4
2L1: for (i = 0; i < 8; i+=4) {
3 S0: a[i] = b[i] * c[i];
4 S1: a[i+1] = b[i+1] * c[i+1];
5 S2: a[i+2] = b[i+2] * c[i+2];
6 S3: a[i+3] = b[i+3] * c[i+3];
7}
Listing 7: Example after partial unrolling
In Listing 6, the loop body denoted by $L$ is depicted in the rectangle. The loop L1 is subjected to a partial unrolling with a factor of 4, as specified by the directive #pragma ACCEL parallel UF=4 hence S0 have to be expanded and replicated four times. The explicit unrolling of the loop is illustrated in Listing 7, where S0, S1, S2 and S3 represent the loop body $L^{{}^{\prime}}$ of the Theorem B.10. The latency of the program in Listing 6 and 7 (which are equivalent) is greater or equal to $\lfloor\frac{TC}{UF}\rfloor$, in this case 2, multiply by the latency of the loop body $L^{{}^{\prime}}$. In this context, $L^{{}^{\prime}}$ comprises statements S0, S1, S2, and S3. These four statements are independent of each other and can be executed concurrently. Therefore, the lower bound on latency for $L^{{}^{\prime}}$ is equivalent to the latency of the multiplication operation.
@@ 6 S3: a[i+3] = b[i+3] * c[i+3];
Consider the scenario of loop $L0$ within Listing 5, which has been unrolled by a factor denoted as $UF\leq TC_{L0}$ where $TC_{L0}$ is the trip count of $L0$. The latency for one iteration of $S0$ is denoted as $Lat^{S0}_{R_{o}}>0$. In the absence of pipelining, the lower bound of the latency of this sub-loop body is: $\lfloor TC_{L0}/UF\rfloor\times Lat^{S0}_{R_{o}}$.
Vitis allows to do a reduction with a tree reduction in logarithmic time with the option “unsafe-math”.
@@
1L1: for (i = 0; i < 8; i++) {
2 c += a[i];
3}
Listing 8: Example of code demonstrating a reduction, where a tree reduction technique can be applied, as depicted in Figure 1.
Figure 1. Illustrating the concurrent execution depicted in Listing 12 through a tree reduction method, accomplished within logarithmic time.
Figure 1 represent the tree reduction of the Listing 12. Due to the reduction, we use a tree reduction in order to increase the parallelism of the reduction, and execute a loop of trip count $n=8$ in $\log_{2}(n)=3$ iterations.
@@ 4.2.3. Loop pipelining
Loop pipelining amounts to overlapping multiple iterations of the loop, so that the next iteration can start prior to the completion of the preceding one. The initiation interval (II) measures in cycles the delay between the start of two consecutive iterations. It is easy to prove our formula template accurately integrates the latency of pipelined loops with the $I$ operator. We compute the minimal II in function of the dependencies of the pipelined loop and the iteration latency of the operations of the statements during the NLP generation. Let $RecMII$ and $ResMII$ be the recurrence constraints and the resource constraints of the pipelined loop, respectively. We have $II\geq\max{(ResMII,RecMII)}$. $RecMII=\max_{i}\lceil\frac{delay(c_{i})}{distance(c_{i})}\rceil$ with $delay(c_{i})$ the total latency in dependency cycle $c_{i}$ and $distance(c_{i})$ the total distance in dependency cycle $c_{i}$. We suppose that $ResMII=1$, as we do not know how the resource will be used by the compiler. Hence, if the loop is a reduction loop then the $II\geq\frac{IL_{reduction}}{1}$ with $IL_{reduction}$ the iteration latency of the operation of reduction. For a kernel like the Listing 9 the $II\geq\lceil\frac{IL_{+}}{2}\rceil$.
@@
Listing 9: Demonstration of a code snippet showcasing a scenario where a loop pipelined with a dependency of distance 2 results in an initiation interval ($II$) that satisfies $II\geq\lceil\frac{IL_{+}}{2}\rceil$.
@@
In Listing 5, loop $L1$ is pipelined. However, loop $L1$ serves is a reduction loop for statement $S2$, meaning that we must await the completion of writing $s[j]$ at iteration $i$ before we can read and write again at iteration $i+1$. Consequently, the initiation interval ($II$) is bounded by or greater than the latency ($IL_{+}$) of the addition operation, which constitutes the reduction operation.
@@ 4.2.7. Program latency lower bound under resource constraints
Given a program and the available resource of DSP $DSP_{avail}$, if $R^{min}_{used}<DSP_{avail}$ the lower bound is valid and the program does not over-utilize the resources.
In Listing 5, statements S0 and S1 entail no DSP usage as they solely serve for initialization purposes. S2 and S3 undergo unrolling N times and operate independently, enabling parallel execution. Particularly, S3 necessitates a tree reduction since it acts as a reduction for loop $L2$.
Assuming 1 DSP for addition and 2 for multiplication (denoted as $DSP_{+}$ and $DSP_{*}$ respectively), the maximum number of computational units utilized by S2 and S3 for both addition and multiplication is N. Therefore, $MCU_{+}^{S2}=N$, $MCU_{*}^{S2}=N$, $MCU_{+}^{S3}=N$, and $MCU_{*}^{S3}=N$. Furthermore, each statement entails only one addition and one multiplication, hence #$\#L_{+}^{S2}=1$, #$\#L_{*}^{S2}=1$, #$\#L_{+}^{S3}=1$, and #$\#L_{*}^{S3}=1$.
Consequently, the minimum number of DSPs utilized by Listing 5 is $N\times(1+2)\times 2$.
@@ 4.2.8. Memory transfer
Given a loop body $L$, the set of array $\mathcal{A}$, an array $a\in\mathcal{A}$, and $LAT^{mem}_{a}$ the latency to transfer the array $a$ from off-chip to on-chip (inputs) and from on-chip to off-chip (outputs). $\forall a\in\mathcal{A},LAT^{mem}_{a}\geq(\mathds{1}_{a\in V_{O}^{L}}+\mathds{1}_{a\in V_{I}^{L}})\times footprint_{a}/max\_burst\_size$. With $\mathds{1}_{cond}=1$ if $cond$ = true else 0.
Within Listing 5, the matrix A is solely read and possesses a footprint of $N\times M\times 32$ bits. Given the FPGA’s limitation to process a maximum of 512 bits per cycle, the latency required for transferring A amounts to $\frac{N\times M}{16}$ cycles.
@@
Given a loop body $L$, the set of array $\mathcal{A}$, the number of cycles to transfer the array $a$ is bounded by $\max_{a\in\mathcal{A}}(\mathds{1}_{a\in V_{O}^{L}}+\mathds{1}_{a\in V_{I}^{L}})\times footprint_{a}/max\_burst\_size$.
In Listing 5, when transferring all arrays before any computation, we can concurrently transfer each array (assuming they are in different DRAM banks). The primary bottleneck arises from the transfer of array $A$, which incurs a latency of $\frac{N\times M}{16}$ cycles. Upon program completion, arrays $s$ and $q$ need to be transferred back to DRAM, with a minimal latency of $\max(\frac{N}{16},\frac{M}{16})$ cycles each. Thus, the total communication latency amounts to $\frac{N\times M}{16}+\max(\frac{N}{16},\frac{M}{16})$ cycles.

### v4

@@ 6.1. A Formal Model for Latency
Our objective is to formulate a lower bound on the latency of a program after HLS. We therefore have put several restrictions: we assume the input program is an affine program, that is the control-flow is statically analyzable; all loops can be recognized and their trip count computed; and all array / memory accesses can be exactly modeled at compile-time. No conditional can occur in the program. While our approach may generalize beyond this specific class, we limit it here to these strict assumptions.
@@
We assume programs are made of affine loops, that are loops with statically computable control-flow, with loop bounds made only of intersection of affine expressions of surrounding loop iterators and program constants. We now assume loop bodies (i.e., statements surrounded by loops) have been translated to a list of statements, with at most a single operation (e.g., +,-,/,*) per statement. Operations are n-ary, that is they take $n\geq 0$ input scalar values as operand, and produce 0 or 1 output scalar value. A memory location can be loaded from (resp. stored to) an address stored in a scalar variable. This is often referred to as straight-line code . This normalization of the loop body facilitates the computation of live-in/live-out data for the code block, and the extraction of the computation graph. Note the region can be represented in Static Single Assignment form, to ensure different storage location for every assignment, facilitating the construction of the operation graph. In addition we require the input program to not contain useless operations which may be removed by the HLS toolchain e.g. by dead code elimination.
@@
  • No HLS optimization shall change the number of operations in the computation: strength reduction, common sub-expression elimination, etc. shall either first be performed in the input program before analysis, or not be performed by the HLS toolchain. The program also does not contain ”useless” operations that may be removed by the compiler with a dead-code elimination pass.
@@
  • We assume that resource (DSP) sharing across different operations executing at the same clock cycle is not possible.
An important term is $SL$, a latency lower bound for a region of straight-line code. To maintain a lower bound on latency by composition, we operate on a representation of (parts of the) program which is both schedule-independent and storage-independent: the operation graph, or CDAG (Hong and Kung, 1981; Elango et al., 2015). Indeed, a lower bound on this representation is necessarily valid under any schedule and storage eventually implemented, and can be used to prove I/O lower bounds on programs (Elango et al., 2015). The CDAG of a straight-line code region is the directed acyclic graph with one node per operation in the code region, connecting all immediate producer and consumer operations directly. Then, we can easily compute the length of its critical path, which represents the minimal set of operations to execute serially.
Definition 6.1 (Straight-line code).
Definition 6.2 (Live-in set).
Definition 6.3 (Live-out set).
@@
Given a straight-line code region $R$ made of a list $L$ of statements $S\in L$, the operation graph OG is the directed graph <$<\{N,V_{I},root,V_{O}\},E>$ such that for every statement $S_{i}$ in the list $L$, there exists a corresponding node $N_{S_{i}}$ in the set of nodes $N$: $\forall S_{i}\in L,~{}N_{S_{i}}\in N$; for each statement $S_{i}$ in the list $L$, the statement produces an output $o_{S_{i}}$ and takes inputs $\vec{i}_{S_{i}}$. $\forall S_{i},\forall i_{k}\in\vec{i}_{S_{i}},e_{i_{k},o}\in E$, $\forall S_{i}:(o_{S_{i}},{\vec{i}}_{S_{i}})\in L,S_{j}:(o_{S_{j}},{\vec{i}}_{S_{j}})\in L$ with $S_{i}\neq S_{j}$ then we have $E_{S_{i},S_{j}}\in E$ iff $o_{Si}\cap{\vec{i}}_{S_{j}}\neq\emptyset$. For every input (resp. output) in $S_{i}$ which is not matched with an output (resp. input) of another $S_{j}$ in $L$, create a node $V_{val}\in V_{I}$ (resp. $V_{O}$) for this input (resp. output) value. If $dim({\vec{i}}_{S_{i}})=0$ then an edge $e_{root,S_{i}}$ is added to $E$.
This formulation ensures that the operation graph effectively represents the flow of data and dependencies between computational statements. By creating edges from inputs to outputs and linking operations with shared data, the graph encodes the structural relationships essential for analyzing or optimizing the execution of straight-line code regions. This representation allows for the straightforward identification of key properties, such as the span or critical path, which are crucial for estimating the latency and area of the code region.
@@ 6.1.1. Latency Lower Bound
We can build a lower bound on the latency of an operation graph:
@@
Given $R_{op}$ a count of available resources of type $op$, for each operation type, let $LO(op)$ be the latency function for operation $op$, with $LO(op)\geq 1$. #$\#L(op)$ denotes the number of operations of type $op$ in $L$. We define $LO(\#OG_{cp}^{L})=\sum_{n\in cp}LO(n)$ the critical path weighted by latency of its operations. The minimal latency of a region $L$ is bounded by
$Lat_{R_{op}}^{L}\geq\max(LO(\#OG_{cp}^{L}),\max_{o\in op}(\lceil LO(o)\times\#L(o)/R_{o}\rceil))$
This theorem provides the building block to our analysis: if reasoning on a straight-line code region, without any loop, then building the operation graph for this region and reasoning on its critical path is sufficient to provide a latency lower bound. As a reminder, all proofs are provided in Appendix A.
1L1: for (i = 0; i < N; i++)
@@ 7 S3: q[i]+=A[i][j]*p[j];
9}
Listing 4: Bicg code: $s=r\times A;q=A\times p$
For example, in Listing 4, if we consider the sub-loop body composed of loops L3 (fully unrolled) and the statements $S2$ and $S3$ as straight-line code regions, we can calculate the critical paths for $S2$ and $S3$ as follows: For $S2$, the critical path is given by: $cp_{S2}=\max(L(+)+L(*),N\times(DSP_{+}+DSP_{*})/R_{o})$ with $DSP_{o}$ the number of DSP for the operation $o$. For $S3$, the critical path is determined by: $cp_{S3}=\max(L(+)\times\log(N)+L(*),N\times(DSP_{+}+DSP_{*})/R_{o})$, considering the possibility of a tree reduction. In this context, the critical path for the entire sub-loop body is the maximum of these two individual critical paths, expressed as $\max(cp_{S2},cp_{S3})$.
@@
6.1.2. Loop Unrolling: full unroll
It follows a simple corollary:
Corollary 6.8 (Equivalence between fully unrolled and straight-line code).
Given a loop nest $l$, if full unrolling is applied to $l$ then the code obtained after full unrolling is a straight-line code as per Def. 6.1.
Consequently, we can bound the latency of a fully unrolled loop nest:
Theorem 6.9 (Minimal latency of a fully unrolled loop nest).
$Lat_{R_{op}}^{l}\geq\max(LO(\#OG_{cp}^{L}),\max_{o\in op}(\lceil L_{o}\times\#L(o)/R_{o}\rceil))$
@@ 6.1.3. Loop Unrolling: partial unroll
Unrolling can be viewed as a two-step transformation: first, strip-mine the loop by the unroll factor, then fully unroll the resulting inner loop. The latency of the resulting sub-program is influenced by how the generated outer loop is implemented. We assume without additional explicit information this unrolled loop will execute in a non-pipelined, non-parallel fashion. Note this bound requires to build the operation graph for the whole loop body. This is straightforward for inner loops and/or fully unrolled loop nests, but impractical if the loop body contains other loops. We therefore define a weaker, but more practical, bound that enables composition:
@@
Consider the scenario of loop $L0$ within Listing 4, which has been unrolled by a factor denoted as $UF\leq TC_{L0}$ where $TC_{L0}$ is the trip count of $L0$. The latency for one iteration of $S0$ is denoted as $Lat^{S0}_{R_{o}}>0$. In the absence of pipelining, the lower bound of the latency of this sub-loop body is: $\lfloor TC_{L0}/UF\rfloor\times Lat^{S0}_{R_{o}}$.
The Vitis HLS compiler enables reductions to be performed using a tree-based reduction algorithm, achieving logarithmic time complexity. This is possible by enabling the unsafe_math_optimizations option, as described in the Xilinx documentation.
@@ 6.1.4. Loop pipelining
Loop pipelining amounts to overlapping multiple iterations of the loop, so that the next iteration can start prior to the completion of the preceding one. The initiation interval (II) measures in cycles the delay between the start of two consecutive iterations. It is easy to prove our formula template accurately integrates the latency of pipelined loops with the $I$ operator. We compute the minimal II in function of the dependencies of the pipelined loop and the iteration latency of the operations of the statements during the NLP generation. Let $RecMII$ and $ResMII$ be the recurrence constraints and the resource constraints of the pipelined loop, respectively. We have $II\geq\max{(ResMII,RecMII)}$. $RecMII=\max_{i}\lceil\frac{delay(c_{i})}{distance(c_{i})}\rceil$ with $delay(c_{i})$ the total latency in dependency cycle $c_{i}$ and $distance(c_{i})$ the total distance in dependency cycle $c_{i}$. We suppose that $ResMII=1$, as we do not know how the resource will be used by the compiler. Hence, if the loop is a reduction loop then the $II\geq\frac{IL_{reduction}}{1}$ with $IL_{reduction}$ the iteration latency of the operation of reduction. For a kernel like the Listing 5 the $II\geq\lceil\frac{IL_{+}}{2}\rceil$.
All dependencies are computed using PolyOpt, and the highest possible value for II is selected based on the dependencies, as defined by the formula.
@@
Listing 5: Demonstration of a code snippet showcasing a scenario where a loop pipelined with a dependency of distance 2 results in an initiation interval ($II$) that satisfies $II\geq\lceil\frac{IL_{+}}{2}\rceil$.
@@ 6.1.8. Program latency lower bound under resource constraints
Given a program and the available resource of DSP $DSP_{avail}$, if $R^{min}_{used}<DSP_{avail}$ the lower bound is valid and the program does not over-utilize the resources. In practice, $MCU_{op}^{s}$ is determined by the unroll factor, which specifies how many times computational units need to be duplicated. However, if the initiation interval (II) of the pipelined loop is strictly greater than one, resource reuse becomes possible within the same statement, reducing the overall number of computational units required.
@@ 6.1.9. Memory transfer
Given a loop body $L$, the set of array $\mathcal{A}$, an array $a\in\mathcal{A}$, and $Lat^{mem}_{a}$ the latency to transfer the array $a$ from off-chip to on-chip (inputs) and from on-chip to off-chip (outputs).
$\left.\begin{cases}\forall a\in\mathcal{A},Lat^{mem}_{a}\geq\frac{footprint_{a}}{max\_burst\_size},&\text{ if $a$ is only read or only write i.e. }(a\in V_{I}^{L}\text{ and }a\notin V_{O}^{L})\\ &\text{ or }(a\notin V_{I}^{L}\text{ and }a\in V_{O}^{L})\\ \forall a\in\mathcal{A},Lat^{mem}_{a}\geq 2\times\frac{footprint_{a}}{max\_burst\_size},&\text{ if $a$ is read and write i.e. }a\in V_{I}^{L}\text{ and }a\in V_{O}^{L}\\ \end{cases}\right.$
@@
Given a loop body $L$, the set of arrays $\mathcal{A}$, the set of loops $\mathcal{L}$, the booleans $loop\_l\_cache\_array\_a$ to indicate if the array $a$ is transferred under the loop $l$, and $footprint_{a}$, the footprint of the array $a$, the latency for communication $Lat^{L}_{communication}$ is bounded by:
$Lat^{L}_{communication}\geq\sum_{l\in\mathcal{L}}\max_{a\in\mathcal{A}}(loop_{l}\_cache\_array\_a\times\frac{footprint_{a}}{max\_burst\_size})$